OpenStudy (sohailiftikhar):

which of the following is greater 70^71 or 71^70

OpenStudy (kropot72):

Take logs of each to find the answer.

OpenStudy (michele_laino):

hint: we can write this: \[\Large \frac{{{{71}^{70}}}}{{{{70}^{71}}}} = {\left( {\frac{{71}}{{70}}} \right)^{70}} \cdot \frac{1}{{70}} = {\left( {1 + \frac{1}{{70}}} \right)^{70}} \cdot \frac{1}{{70}}\]

OpenStudy (anonymous):

To me, \(71^{70}\) is greater.

OpenStudy (sohailiftikhar):

no 70^71 is greater